# Frequency-Guided Masking for Enhanced Vision Self-Supervised Learning

### v2

@@ Appendix D Prediction Visualization
![[Uncaptioned image]](extracted/6275800/images/1/org.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq_50.png)
![[Uncaptioned image]](extracted/6275800/images/1/org_50.png)
![[Uncaptioned image]](extracted/6275800/images/1/_out_50.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/org_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/_out_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq_10.png)
![[Uncaptioned image]](extracted/6275800/images/1/org_10.png)
![[Uncaptioned image]](extracted/6275800/images/1/_out_10.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/org_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/_out_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq_5.png)
![[Uncaptioned image]](extracted/6275800/images/1/org_5.png)
![[Uncaptioned image]](extracted/6275800/images/1/_out_5.png)
![[Uncaptioned image]](extracted/6275800/images/1/freq_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/org_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/1/_out_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/org.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq_50.png)
![[Uncaptioned image]](extracted/6275800/images/2/org_50.png)
![[Uncaptioned image]](extracted/6275800/images/2/_out_50.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/org_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/_out_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq_10.png)
![[Uncaptioned image]](extracted/6275800/images/2/org_10.png)
![[Uncaptioned image]](extracted/6275800/images/2/_out_10.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/org_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/_out_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq_5.png)
![[Uncaptioned image]](extracted/6275800/images/2/org_5.png)
![[Uncaptioned image]](extracted/6275800/images/2/_out_5.png)
![[Uncaptioned image]](extracted/6275800/images/2/freq_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/org_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/2/_out_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/org.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq_50.png)
![[Uncaptioned image]](extracted/6275800/images/3/org_50.png)
![[Uncaptioned image]](extracted/6275800/images/3/_out_50.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/org_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/_out_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq_10.png)
![[Uncaptioned image]](extracted/6275800/images/3/org_10.png)
![[Uncaptioned image]](extracted/6275800/images/3/_out_10.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/org_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/_out_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq_5.png)
![[Uncaptioned image]](extracted/6275800/images/3/org_5.png)
![[Uncaptioned image]](extracted/6275800/images/3/_out_5.png)
![[Uncaptioned image]](extracted/6275800/images/3/freq_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/org_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/3/_out_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/org.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq_50.png)
![[Uncaptioned image]](extracted/6275800/images/4/org_50.png)
![[Uncaptioned image]](extracted/6275800/images/4/_out_50.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/org_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/_out_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq_10.png)
![[Uncaptioned image]](extracted/6275800/images/4/org_10.png)
![[Uncaptioned image]](extracted/6275800/images/4/_out_10.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/org_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/_out_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq_5.png)
![[Uncaptioned image]](extracted/6275800/images/4/org_5.png)
![[Uncaptioned image]](extracted/6275800/images/4/_out_5.png)
![[Uncaptioned image]](extracted/6275800/images/4/freq_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/org_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/4/_out_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/org.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq_50.png)
![[Uncaptioned image]](extracted/6275800/images/5/org_50.png)
![[Uncaptioned image]](extracted/6275800/images/5/_out_50.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/org_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/_out_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq_10.png)
![[Uncaptioned image]](extracted/6275800/images/5/org_10.png)
![[Uncaptioned image]](extracted/6275800/images/5/_out_10.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/org_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/_out_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq_5.png)
![[Uncaptioned image]](extracted/6275800/images/5/org_5.png)
![[Uncaptioned image]](extracted/6275800/images/5/_out_5.png)
![[Uncaptioned image]](extracted/6275800/images/5/freq_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/org_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/5/_out_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/org.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq_50.png)
![[Uncaptioned image]](extracted/6275800/images/6/org_50.png)
![[Uncaptioned image]](extracted/6275800/images/6/_out_50.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/org_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/_out_50_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq_10.png)
![[Uncaptioned image]](extracted/6275800/images/6/org_10.png)
![[Uncaptioned image]](extracted/6275800/images/6/_out_10.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/org_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/_out_10_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq_5.png)
![[Uncaptioned image]](extracted/6275800/images/6/org_5.png)
![[Uncaptioned image]](extracted/6275800/images/6/_out_5.png)
![[Uncaptioned image]](extracted/6275800/images/6/freq_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/org_5_m.png)
![[Uncaptioned image]](extracted/6275800/images/6/_out_5_m.png)

### v3

@@ Appendix D Prediction Visualization
![[Uncaptioned image]](extracted/6319251/images/1/org.png)
![[Uncaptioned image]](extracted/6319251/images/1/freq.png)
![[Uncaptioned image]](extracted/6319251/images/1/freq_50.png)
![[Uncaptioned image]](extracted/6319251/images/1/org_50.png)
![[Uncaptioned image]](extracted/6319251/images/1/_out_50.png)
![[Uncaptioned image]](extracted/6319251/images/1/freq_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/1/org_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/1/_out_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/1/freq_10.png)
![[Uncaptioned image]](extracted/6319251/images/1/org_10.png)
![[Uncaptioned image]](extracted/6319251/images/1/_out_10.png)
![[Uncaptioned image]](extracted/6319251/images/1/freq_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/1/org_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/1/_out_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/1/freq_5.png)
![[Uncaptioned image]](extracted/6319251/images/1/org_5.png)
![[Uncaptioned image]](extracted/6319251/images/1/_out_5.png)
![[Uncaptioned image]](extracted/6319251/images/1/freq_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/1/org_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/1/_out_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/2/org.png)
![[Uncaptioned image]](extracted/6319251/images/2/freq.png)
![[Uncaptioned image]](extracted/6319251/images/2/freq_50.png)
![[Uncaptioned image]](extracted/6319251/images/2/org_50.png)
![[Uncaptioned image]](extracted/6319251/images/2/_out_50.png)
![[Uncaptioned image]](extracted/6319251/images/2/freq_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/2/org_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/2/_out_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/2/freq_10.png)
![[Uncaptioned image]](extracted/6319251/images/2/org_10.png)
![[Uncaptioned image]](extracted/6319251/images/2/_out_10.png)
![[Uncaptioned image]](extracted/6319251/images/2/freq_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/2/org_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/2/_out_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/2/freq_5.png)
![[Uncaptioned image]](extracted/6319251/images/2/org_5.png)
![[Uncaptioned image]](extracted/6319251/images/2/_out_5.png)
![[Uncaptioned image]](extracted/6319251/images/2/freq_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/2/org_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/2/_out_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/3/org.png)
![[Uncaptioned image]](extracted/6319251/images/3/freq.png)
![[Uncaptioned image]](extracted/6319251/images/3/freq_50.png)
![[Uncaptioned image]](extracted/6319251/images/3/org_50.png)
![[Uncaptioned image]](extracted/6319251/images/3/_out_50.png)
![[Uncaptioned image]](extracted/6319251/images/3/freq_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/3/org_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/3/_out_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/3/freq_10.png)
![[Uncaptioned image]](extracted/6319251/images/3/org_10.png)
![[Uncaptioned image]](extracted/6319251/images/3/_out_10.png)
![[Uncaptioned image]](extracted/6319251/images/3/freq_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/3/org_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/3/_out_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/3/freq_5.png)
![[Uncaptioned image]](extracted/6319251/images/3/org_5.png)
![[Uncaptioned image]](extracted/6319251/images/3/_out_5.png)
![[Uncaptioned image]](extracted/6319251/images/3/freq_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/3/org_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/3/_out_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/4/org.png)
![[Uncaptioned image]](extracted/6319251/images/4/freq.png)
![[Uncaptioned image]](extracted/6319251/images/4/freq_50.png)
![[Uncaptioned image]](extracted/6319251/images/4/org_50.png)
![[Uncaptioned image]](extracted/6319251/images/4/_out_50.png)
![[Uncaptioned image]](extracted/6319251/images/4/freq_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/4/org_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/4/_out_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/4/freq_10.png)
![[Uncaptioned image]](extracted/6319251/images/4/org_10.png)
![[Uncaptioned image]](extracted/6319251/images/4/_out_10.png)
![[Uncaptioned image]](extracted/6319251/images/4/freq_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/4/org_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/4/_out_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/4/freq_5.png)
![[Uncaptioned image]](extracted/6319251/images/4/org_5.png)
![[Uncaptioned image]](extracted/6319251/images/4/_out_5.png)
![[Uncaptioned image]](extracted/6319251/images/4/freq_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/4/org_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/4/_out_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/5/org.png)
![[Uncaptioned image]](extracted/6319251/images/5/freq.png)
![[Uncaptioned image]](extracted/6319251/images/5/freq_50.png)
![[Uncaptioned image]](extracted/6319251/images/5/org_50.png)
![[Uncaptioned image]](extracted/6319251/images/5/_out_50.png)
![[Uncaptioned image]](extracted/6319251/images/5/freq_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/5/org_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/5/_out_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/5/freq_10.png)
![[Uncaptioned image]](extracted/6319251/images/5/org_10.png)
![[Uncaptioned image]](extracted/6319251/images/5/_out_10.png)
![[Uncaptioned image]](extracted/6319251/images/5/freq_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/5/org_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/5/_out_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/5/freq_5.png)
![[Uncaptioned image]](extracted/6319251/images/5/org_5.png)
![[Uncaptioned image]](extracted/6319251/images/5/_out_5.png)
![[Uncaptioned image]](extracted/6319251/images/5/freq_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/5/org_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/5/_out_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/6/org.png)
![[Uncaptioned image]](extracted/6319251/images/6/freq.png)
![[Uncaptioned image]](extracted/6319251/images/6/freq_50.png)
![[Uncaptioned image]](extracted/6319251/images/6/org_50.png)
![[Uncaptioned image]](extracted/6319251/images/6/_out_50.png)
![[Uncaptioned image]](extracted/6319251/images/6/freq_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/6/org_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/6/_out_50_m.png)
![[Uncaptioned image]](extracted/6319251/images/6/freq_10.png)
![[Uncaptioned image]](extracted/6319251/images/6/org_10.png)
![[Uncaptioned image]](extracted/6319251/images/6/_out_10.png)
![[Uncaptioned image]](extracted/6319251/images/6/freq_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/6/org_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/6/_out_10_m.png)
![[Uncaptioned image]](extracted/6319251/images/6/freq_5.png)
![[Uncaptioned image]](extracted/6319251/images/6/org_5.png)
![[Uncaptioned image]](extracted/6319251/images/6/_out_5.png)
![[Uncaptioned image]](extracted/6319251/images/6/freq_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/6/org_5_m.png)
![[Uncaptioned image]](extracted/6319251/images/6/_out_5_m.png)